f(x)=3x+1 and g(x)=x+2 :
find (f+g(x)
calculate (f+g)(1)

Answer by rothauserc(4718)

(f+g(x)) = f(x) + g(x) = 3x+1 + x+2 = 4x+3
(f+g(1)) = 4*1 +3 = 7